OpenStudy (anonymous):

What is a possible first step to eliminate the variable x in the following system? 2x – 7y = 3 4x + 5y = –2

Directrix (directrix):

Multiply each member of this equation by (-2) and post the result here. Equation: 2x – 7y = 3

OpenStudy (anonymous):

Top one by -2 :) -4x + 14y = -6 4x + 5y = -2 19y = -8 and you can work it from there :)

Directrix (directrix):

@ryan.g.campbell.17 ---> This question was for the Asker and not for your deus ex machina. Read the Code of Conduct. @raebaby420 Yes. The objective in the Addition/Subtraction method of solving simultanteous linear equations is to multiply by a number that will cause the coefficient of the variable you wish to eliminate to add out.

OpenStudy (anonymous):

Ok.. so i think i understand it

OpenStudy (anonymous):

Soo i just gotta single out my coefficient?

Directrix (directrix):

@raebaby420 Let's try another one. Same instructions. Answer? 3x + 4y = 7 3x - 4y = 10

OpenStudy (anonymous):

umm haha i have no clue.

OpenStudy (anonymous):

whats the first step i have to complete?

Directrix (directrix):

Question: What added to 3x gives 0? (3x) + ? = 0 ? =